# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multi-modal Foundation Models

### v3

@@ 5.1 Dataset
5.2 Generative Models under Inductive Biases
Model Method BLEU ROUGE-L METEOR SPICE Accuracy
DDPM LatentExplainer 25.1 39.1 36.9 26.0 76.3
w/o Inductive Bias Prompts 4.0 15.2 12.9 8.4 32.4
w/o Uncertainty Quantification 16.8 32.0 29.0 19.8 70.0
w/o IB and w/o UQ 2.9 13.8 11.8 7.1 32.2
$\beta$-TCVAE LatentExplainer 22.6 47.3 35.6 24.9 82.4
w/o Inductive Bias Prompts 3.8 27.5 20.7 10.5 52.9
w/o Uncertainty Quantification 18.7 41.0 33.8 22.6 54.9
w/o IB and w/o UQ 3.5 27.5 19.4 10.7 47.1
Table 5: Ablation study of the proposed LatentExpainer on inductive bias prompts and the selection of explanations based on the uncertainty for the disentanglement bias.
Model Dataset BLEU ROUGE-L METEOR SPICE Accuracy
CSVAE LatentExplainer 25.8 41.5 35.2 28.5 81.0
w/o Inductive Bias Prompts 8.7 27.4 21.2 12.2 54.6
w/o Uncertainty Quantification 16.2 35.6 27.6 22.8 66.7
w/o IB and w/o UQ 5.8 27.3 20.2 10.6 52.4
Table 6: Ablation study of the proposed LatentExpainer on inductive bias prompts and the selection of explanations based on the uncertainty for the combination bias.
Model Dataset BLEU ROUGE-L METEOR SPICE Accuracy
Stable Diffusion LatentExplainer 16.7 36.2 30.0 23.0 89.3
w/o Inductive Bias Prompts 8.9 27.6 24.0 15.5 63.1
w/o Uncertainty Quantification 14.9 32.8 27.0 22.0 88.4
w/o IB and w/o UQ 7.1 25.7 21.8 14.7 55.6
CSVAE LatentExplainer 23.7 39.1 32.2 20.3 75.6
w/o Inductive Bias Prompts 11.0 29.0 22.8 12.2 48.9
w/o Uncertainty Quantification 11.8 29.5 22.1 14.2 57.8
w/o IB and w/o UQ 6.5 23.0 17.0 8.5 44.4
Table 7: Ablation study of the proposed LatentExpainer on inductive bias prompts and the selection of explanations based on the uncertainty for the conditional bias.
5.3 Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), SPICE Anderson et al. (2016) as the automated metrics and human annotated accuracy to assess the explanations of latent variables. In Table 2, the explanations generated with the inductive bias prompts consistently outperform their counterparts without inductive bias prompts across all datasets and all metrics. Similar results can be found in Table 3 and Table 4. Their consistent results demonstrate that inductive bias is very important and necessary when explaining the latent representations of generative models. They also show that our framework can effectively design prompts for different inductive biases in generative models to improve the accuracy of latent explanations.
@@ 5.4 Qualitative Evaluation
To analyze the explanations for DDPM under disentanglement bias, we manipulate the latent representation along each latent direction and compare it with the one that first traverses along another latent direction and then traverses along the same latent direction. In Figure 3, for each latent direction, we pass two image sequences perturbed along that latent direction and the inductive bias prompt based on the disentanglement formula to MLLM to obtain the common latent explanation through the uncertainty quantification. In comparison, the explanations generated without the inductive bias prompt as shown in Figure 4 tend to show "no clear explanation" or explanations that align with one image sequence but do not reflect the common pattern in both image sequences. The addition of the inductive bias prompts can assist with ruling out the variation effects of other latent variables.
We also qualitatively evaluate the explanations for CSVAE under combination bias. We transverse each latent dimension and compare it with the one that first changes another latent variable in another group and then traverses the same latent variable, as well as the ones that first change another latent variable in the same group and then traverse the same latent variable. As Figure 10 depicts, our model can clearly show the explanations of latent variable $z_{1},z_{2},z_{3}$ as the color of the ground, the background color, and the shape of the object. The effect of removing inductive bias prompts is similar to the disentanglement bias.
In Figure 11, we provide the "young appearance" prompt to Stable Diffusion under conditional bias, and all three top latent directions are associated with the prompt. Compared with the one without the inductive bias prompts in Figure 12, the addition of inductive bias prompts can better identify the relation with the property of interest in the explanations as we can see the suggesting meaning with the property in addition to the description of the pattern in both $v_{1}$ and $v_{2}$. More evaluation results can be found in Appendix C.
5.5 Ablation Study
To understand the impact of various components in our proposed LatentExplainer, we conducted a comprehensive ablation study across different models (DDPM, $\beta$-TCVAE, Stable Diffusion, and CSVAE) and inductive biases (disentanglement, combination, and conditional). Specifically, we evaluated the removal of inductive bias prompts (IB), uncertainty quantification (UQ), and both components. The results are summarized in Tables 5, Tables 6, and Tables 7. Removing inductive bias prompts leads to a substantial drop in all generative models, indicating the importance of inductive bias in generating high-quality explanations. Likewise, the removal of uncertainty quantification also results in decreased performance in all generative models. Overall, the ablation studies confirm the necessity of both inductive bias prompts and uncertainty quantification in our LatentExplainer framework, demonstrating their significant contributions to improving explanatory performance across various generative models.
@@ 6 Conclusion
In this paper, we introduced LatentExplainer, a framework designed to generate semantically meaningful explanations of latent variables in deep generative models. Our experiments quantitatively revealed that the inclusion of inductive bias prompts significantly enhances the quality of explanations generated by LatentExplainer. In qualitative evaluations, LatentExplainer provided clear explanations for the latent variables, effectively reflecting superior performance when inductive bias prompts were utilized. The ablation studies further confirmed the necessity of inductive bias prompts and uncertainty quantification in producing accurate and consistent explanations.
@@ 7 Limitation
Appendix A Human Annotaions
The ground-truth annotations of explanations are performed by four different annotators from the United States and China and the results are then aggregated to calculate the automated evaluation metrics. To obtain the accuracy, the annotators have to select one out of four choices(yes, weak yes, weak no, and no) as a response for whether the explanation justifies the pattern of change in the image sequence. The four levels of rating are then mapped to the scores of 1, 2/3, 1/3, and 0 respectively. The scores are then averaged among data samples in the datasets to have the accuracy values, following the evaluation process in  Sammani et al. (2022); Chang et al. (2023).
Appendix B Selection of Threshold of the Certainty Scores
We follow our previous work Zhu et al. (2024) to find the threshold, we denote the true label of the interpretability of each latent variable as 1 if at least two annotators out of three can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of each response is its predicted probability. We can then compute its AUC (area under the ROC curve) based on different thresholds and choose the one with the highest AUC as our threshold $\varepsilon$. Our final output $\mathcal{T}$ is the response that has the highest mean pairwise cosine similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, $\mathcal{T}$ will be "no clear explanation".
Appendix C More Qualitative Evaluation Results
@@
We provide more qualitative evaluation results: (1) for the disentanglement bias in Figure 5, and Figure 6; (2) for the combination bias in Figure 7; (3) for the conditional bias in Figure 8 and Figure 9.
@@
Appendix D Visualization of the Generated Explanations

### v4

@@ 5.5 Qualitative Evaluation
To analyze the explanations for DDPM under disentanglement bias, we manipulate the latent representation along a latent direction and compare it with the one that first traverses along another latent direction and then traverses along the same latent direction. The disentangled latent variable would be invariant with respect to the variations in another latent dimension. In Figure 3, for each latent direction, we pass two image sequences and an inductive bias prompt based on the disentanglement formula to the MLLM to obtain a common latent explanation. In comparison, the explanations generated without the inductive bias prompt as shown in Figure 4 tend to show "no clear explanation" or wrong explanations as they only align with one image sequence but do not reflect the common pattern in both image sequences. The addition of the inductive bias prompts can assist with ruling out the variation effects of other latent variables to capture the actual meaning.
We also qualitatively evaluate the explanations for CSVAE under combination bias. We transverse a latent variable and compare it with the one that first traverses another latent variable in another group and then traverses the same latent variable, which is similar to the disentanglement bias. We then compare with the one that first transverses another latent variable in the same group and then traverses the same latent variable. As Figure 10 depicts, our model can clearly show the explanations of latent variable $z_{1},z_{2},z_{3}$ as the color of the ground, the background color, and the shape of the object. The effect of removing inductive bias prompts leads to no clear explanation or wrong explanations like the disentanglement bias.
In Figure 11, we provide the "young appearance" prompt to Stable Diffusion under conditional bias, and the explanations of all three top latent directions reflect the meaning of youth. The addition of inductive bias prompts can better identify the relation with the property of interest to capture the actual meaning of latent variables. In comparison, the one without the inductive bias prompts in Figure 12, cannot find clear explanations or simply describe the characteristics in the image sequence, lacking an abstract generalization. More qualitative evaluation results can be found in Appendix D.
@@ 6 Conclusion
In this paper, we introduced LatentExplainer, a framework designed to generate semantically meaningful explanations of latent variables in deep generative models. Our experiments quantitatively revealed that the inclusion of inductive bias prompts significantly enhances the quality of explanations generated by LatentExplainer. In qualitative evaluations, LatentExplainer provided clear explanations for the latent variables, effectively reflecting superior performance when inductive bias prompts were utilized. The quantitative analysis further confirmed the necessity of inductive bias prompts and uncertainty quantification in producing accurate and consistent explanations.
@@ 7 Limitation
Appendix A Latent Variable Perturbation
For VAE models, every time we target at each latent variable. The targeted latent variable is changed from -3 to 3, which is the 3 times the standard deviation of the standard Gaussian distribution of a latent variable, with a step size of 1.5. All the other latent variables are unchanged. For diffusion models, we follow $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{v})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is the strength. We set the value of $\gamma$ to 0.1, 0.2, 0.3, 0.4, and 0.5 for DDPM, and to 1, 2, 3, 4, and 5 for Stable Diffusion.
Appendix B Uncertainty-aware Implementations
We denote the true label of the interpretability for the i-th latent variable $y_{i}$ as 1 if at least two of the three annotators can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for the i-th latent variable is $\bar{s}_{i}$. We adopt the Jaccard Index to measure similarity between the predicted label and the true label,
$\begin{split}&E(\varepsilon)=\texttt{Jaccard}(f(\bar{s}_{i},\varepsilon),y_{i}),\text{ where}\\ &f(\bar{s}_{i},\varepsilon)=\mathbbm{1}(\bar{s}_{i}\geq\varepsilon).\end{split}$
The threshold is then selected as the one with the maximum similarity with the true label,
$\varepsilon^{*}=\arg\max_{\varepsilon}E(\varepsilon).$
By solving this equation across all datasets, we find the threshold $\varepsilon=0.2617$. Since $\varepsilon$ is solved across all datasets, it reduces the dependence on any specific dataset and can be applied to new data. Our final output $\mathcal{T}$ is the response that has the highest mean pairwise cosine similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, $\mathcal{T}$ will be "no clear explanation".
Appendix C An example of automatically generating prompts from an inductive bias formula
Example: Given a formula $\mathcal{F}$ which is “$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),\forall t_{1}\neq t_{2}"$, generate a prompt that reflects the inductive bias.
Solution:
Step 1: Extracting mathematical symbols: Extracting all mathematical symbols from the given formula “$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),\forall t_{1}\neq t_{2}"$. Symbols extracted are $\mathcal{S}=\left[\right.$$p\left(z_{i}\mid\cdot\right)"$, “$\left.t=t_{1},t=t_{2},\forall t_{1}\neq t_{2}",``\neq"\right]$.
[…]
Step 3: Generating a prompt for the given formula $\mathcal{F}$ : We can generate a prompt for the given formula $\mathcal{F}$ using a pretrained LLM $\pi_{\theta}$ with the following prompt that consists of the few-shot examples $\mathcal{H}$ of the three inductive biases, $\mathcal{F}$, and the gathered $\mathtt{semantics}$:
Formula:
$p\left(z_{i}\mid z_{i^{\prime}}=\alpha\right)=p\left(z_{i}\mid z_{i^{\prime}}=\beta\right),\forall i\neq i^{\prime},\alpha\neq\beta$.
Prompt: These two rows of images show the same pattern of change despite other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation".
Formula:
$p\left(z_{i}\mid z_{j}=\alpha\right)=p\left(z_{i}\mid z_{j}=\beta\right),\forall z_{i}\in G,z_{j}\in G^{\prime},G\neq G^{\prime},\alpha\neq\beta$.
$p\left(z_{i}\mid z_{i^{\prime}}=\alpha\right)\neq p\left(z_{i}\mid z_{i^{\prime}}=\beta\right),\forall z_{i},z_{i}^{\prime}\in G,i\neq i^{\prime},\alpha\neq\beta.$
Prompt: The pattern of change is associated with a group. The first two rows of images show the same pattern of change despite other variations in another group. The pattern of change in the last two rows of images should change given other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation".
Formula:
$p\left(z_{i}\mid p_{k}=\alpha\right)\neq p\left(z_{i}\mid p_{k}=\beta\right),\forall z_{i}\in G_{k},\alpha\neq\beta.$
$p\left(z_{j}\mid p_{k}=\alpha\right)=p\left(z_{j}\mid p_{k}=\beta\right),\forall z_{j}\notin G_{k},\alpha\neq\beta.$
[…]
Formula:
$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),\forall t_{1}\neq t_{2}$.
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$t=t_{1},t=t_{2},\forall t_{1}\neq t_{2}"$ is “different timesteps"""", “"absent"$\neq$"≠ " is “change"""". Please give a prompt in plain language.
Appendix D More Qualitative Evaluation Results
@@
We provide more qualitative evaluation results: (1) the disentanglement bias: DDPM in Figure 5, and $\beta$-TCVAE in Figure 6; (2) the combination bias: CAVAE in Figure 7; (3) the conditional bias: Stable Diffusion in Figure 8 and CSVAE in Figure 9.
@@
Appendix E Visualization of the Generated Explanations
@@
Model Method AFHQ LSUN-Church CelebA-HQ
[…]
DDPM (Disentanglement Bias) GPT-4o Baseline 3.08 14.83 14.22 8.78 87.55 -3.39 5.52 23.06 16.85 11.23 88.47 -3.41 0.11 3.54 4.44 1.22 85.97 -3.16
+ LatentExplainer w/o IB 2.14 11.49 12.01 7.82 87.17 -3.41 8.78 27.01 20.15 14.08 88.85 -3.37 0.19 4.17 4.34 1.80 85.88 -3.16
+ LatentExplainer w/o UQ 16.12 31.57 31.74 21.70 89.99 -3.12 21.19 37.03 29.40 23.65 90.49 -3.19 13.11 27.38 25.97 14.03 89.15 -2.92
+ LatentExplainer 25.91 37.84 41.52 29.87 91.48 -2.97 30.92 44.21 36.39 29.23 91.80 -3.06 18.49 35.27 32.76 18.81 90.30 -2.90
Stable Diffusion (Conditional Bias) GPT-4o Baseline 7.61 26.32 21.52 17.19 90.04 -3.32 7.80 26.73 23.40 14.28 89.44 -3.46 5.79 23.89 20.44 12.68 89.63 -3.04
+ LatentExplainer w/o IB 9.12 26.59 22.46 16.28 89.77 -3.29 10.07 28.31 25.28 15.85 89.92 -3.42 5.87 24.06 21.09 12.17 89.15 -3.05
+ LatentExplainer w/o UQ 12.36 27.73 22.64 20.01 90.03 -3.18 17.19 35.33 30.50 21.76 90.73 -3.31 15.01 35.43 27.96 23.99 91.07 -2.87
+ LatentExplainer 13.27 29.98 23.56 20.08 90.17 -3.19 18.33 38.14 32.46 25.01 90.99 -3.23 18.50 40.85 33.90 23.90 91.75 -2.82
Table 5: Quantitative results for diffusion models across datasets. B represents BLEU@4, R represents ROUGE-L, M represents METEOR, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
Model Method 3DShapes CelebA-HQ dSprites
[…]
$\beta$-TCVAE (Disentanglement Bias) GPT-4o Baseline 5.51 29.77 19.12 7.92 89.18 -3.13 11.09 37.14 28.80 17.24 90.55 -3.07 0.00 16.19 10.38 7.04 87.17 -3.22
+ LatentExplainer w/o IB 5.41 31.66 22.28 10.07 89.44 -3.13 6.14 32.73 25.33 14.04 89.92 -3.11 0.00 18.73 14.42 7.54 87.42 -3.20
+ LatentExplainer w/o UQ 16.99 37.37 30.72 22.89 90.83 -2.80 21.95 48.84 39.70 22.60 92.09 -2.88 17.16 37.19 31.11 22.40 90.18 -2.89
+ LatentExplainer 25.40 49.06 32.72 22.78 91.90 -2.75 29.93 55.68 44.32 30.17 93.55 -2.77 12.55 37.30 29.86 21.69 90.28 -2.87
CSVAE (Combination Bias) GPT-4o Baseline 6.93 25.07 19.04 9.70 89.42 -2.87 10.44 37.42 28.51 15.06 90.53 -3.01 0.00 21.61 13.17 6.99 87.19 -3.17
+ LatentExplainer w/o IB 14.22 28.57 25.12 11.79 90.02 -2.85 11.96 37.02 28.51 17.49 90.50 -3.01 0.00 16.36 10.12 7.26 86.96 -3.24
+ LatentExplainer w/o UQ 34.55 39.62 33.42 30.30 90.83 -2.62 14.08 38.68 30.13 22.49 91.11 -2.87 0.00 28.49 19.19 15.65 89.78 -2.91
+ LatentExplainer 36.18 43.58 36.60 36.75 91.72 -2.52 25.34 45.18 39.82 26.96 92.09 -2.81 16.03 35.85 29.23 21.82 90.05 -2.90
CSVAE (Conditional Bias) GPT-4o Baseline 10.97 30.55 23.90 8.83 90.10 -3.02 8.46 28.78 21.93 15.78 89.62 -2.98 0.00 9.62 5.27 1.03 85.74 -3.34
+ LatentExplainer w/o IB 19.28 39.90 35.47 15.74 90.87 -2.88 8.35 28.75 18.45 12.11 89.60 -2.97 0.00 11.38 9.35 5.26 86.47 -3.29
+ LatentExplainer w/o UQ 16.73 32.28 25.17 16.19 89.84 -2.86 13.35 37.20 25.40 20.05 89.89 -2.92 5.39 19.20 15.65 6.36 86.94 -3.26
+ LatentExplainer 25.71 40.00 31.59 20.67 90.88 -2.79 24.21 42.88 36.19 19.61 91.21 -2.79 21.17 34.99 28.71 20.71 89.49 -3.08
Table 6: Quantitative results for VAE models across datasets. B represents BLEU@4, R represents ROUGE-L, M represents METEOR, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.